OpenStudy (precal):

ok need help-yes I should also, be able to do this 6th grade math problem.......need a double check Thanks in advance

OpenStudy (precal):

Hector is buying 6 equally prices packages of frozen vegetables. He also, uses a coupon for 1.50 off the total cost. He pays with a 20 bill and gets 3.20 in change. Which method can be used to find the regualr price of each package of frozen vegetables in dollars?

OpenStudy (precal):

20-1.5-3.2=15.3/6=2.55

OpenStudy (precal):

but wait don't forget the discount

OpenStudy (precal):

.25 each one

OpenStudy (precal):

2.55+.25=2.8 each

OpenStudy (precal):

well now here are my lovely choices 1) divide the difference from 20 and 3.20 by 6 and then subtract 1.50 from this quotient

OpenStudy (precal):

2)multiply the difference from 20 and 3.20 by 6 and then add 1.50 to this product

OpenStudy (precal):

3)subtract 1.50 from the difference between 20 and 3.20 and then divide this difference by 6

OpenStudy (precal):

4) add 1.50 to the difference between 20 and 3.20 and then divide this sum by 6

OpenStudy (precal):

ok those are my 4 choices

OpenStudy (precal):

not one because 20-3.2=16.8/6=2.8-1.5=1.3

OpenStudy (anonymous):

Are equations and variables not allowed in 6th grade? Sounds tough...

OpenStudy (precal):

not two b/c 20-3.2=16.8x6=100.8=1.5=102.3

OpenStudy (precal):

I have no idea but I don't think the choices are correct unless I don't know how to read or it could have a typo.....just need a second pair of eyes to make sure I am not overlooking anything

OpenStudy (anonymous):

\[ 6n-1.50=20.00-3.20 \]

OpenStudy (precal):

third choice is 20-3.2-1.5=15.3/6=2.55 but that is the price of each one with the discount......hmmmmmmm

OpenStudy (precal):